Ashley is 64 years old and Stephanie is 18 years old. How many years will it take until Ashley is only 3 times as old as Stephanie?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52786)

.
64 + x = 3*(18+x),   where x is the value under the question.


64 + x = 54 + 3x

64 - 54 = 3x - x


10 = 2x  ====>  x = 10/2 = 5 years.


Answer.  In 5 years.


Check.   In 5 years they will be  69 and 23 years old.   69 = 23*3.   ! Correct !


Answer by greenestamps(13200)



When Ashley is 3 times as old as Stephanie, their ages can be represented as 3x and x; the difference in their ages will be 2x.

The difference in their ages is now, and always will be, 46 years.

2x = 46 --> x = 23; so Ashley will be 3 times as old as Stephanie when Stephanie is 23 and Ashley is 69.

Since Ashley is now 64, that will be 5 years from now.